Consider two solids, one (A) with a density of 2.79 g/mL, and the other (B) with a density of 7.91 g/mL.

a) If you mix 1.00 mL of each of the two substances together, what is the density of the mixture?............g/ml


b) If you mix 1.00 g of each of the two substances together, what is the density of the mixture?..........g/ml



1
Expert's answer
2021-01-25T04:19:55-0500

a)

Masses of each sample are:

"m_A=D_A\times{V_A}=2.79g/mL\times1.00mL=2.79g"

"m_B=D_B\times{V_B}=7.91g/mL\times1.00mL=7.91g"

Therefore, the density of the mixture is:

"D_{AB}=\frac{m_A+m_B}{V_A+V_B}=\frac{2.79g+7.91g}{1.00mL+1.00mL}=5.35g/mL"


Answer: 5.35 g/mL


b)

Volumes of each sample are:

"V_A=\frac{m_A}{D_A}=\frac{1.00g}{2.79g/mL}=0.3584mL"

"V_B=\frac{m_B}{D_B}=\frac{1.00g}{7.91g/mL}=0.1264mL"

Therefore, the density of the mixture is:

"D_{AB}=\frac{m_A+m_B}{V_A+V_B}=\frac{1.00g+1.00g}{0.3584mL+0.1264mL}=4.13g/mL"


Answer: 4.13 g/mL